40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 108 of 341  Not logged in ELOG logo
ID Date Author Type Category Subject
  11739   Mon Nov 9 09:27:44 2015 SteveUpdatePSLMC is not happy

I just turned off the PSL enclousure lights.
 

Attachment 1: ioo8d.png
ioo8d.png
  11738   Fri Nov 6 15:56:00 2015 gautamUpdateLSCFSR and linewidth measurements with phase tracker

Summary:

I performed a preliminary calibration of the X and Y phase trackers, and found that the slopes of a linear fit of phase tracker output as a function of driven frequency (as measured with digital frequency counter) are 0.7886 +/- 0.0016 and 0.9630 +/- 0.0012 respectively (see Attachments #1 and #2). Based on this, the EPICS calibration constants have been updated. The data used for calibration has also been uploaded (Attachment #4).

Details:

I found that by adopting the approach I suggested as a fix in elog 11736, and setting a gate time of 1second, I could eliminate the systematic bias in measured frequency I had been seeing, the origin of which is also discussed in elog 11736. This was verified by using a digital oscillator to supply the input to the frequency counting block, and verifying that I could recover the driving frequency without any systematic bias. Therefore, I used this as a measure of the driving frequency independent of the front panel display of the Fluke 6061A. 

The actual calibration was done as follows:

  1. Close PSL green and end green shutters. Turned off the power to the green transmission PDs on the PSL table and disconnected the couplers from their outputs.
  2. Connected the output of the Fluke 6061A RF signal generator to a splitter, and to the inputs of the couplers for the X and Y signal chains.
  3. Adjusted the amplitude of the RF signal output until the Q readout of the rotated X and Y outputs were between 1000 and 3000. The final value used was -17dBm. As a qualitative check, I also looked at the beat signal on the spectrum analyzer in the control room and judged the peak height to be roughly the same as that seen when a real beat note was being measured. The phase tracker gains after setting the UGF were ~83 and 40 for the X and Y arms respectively.
  4. Step through the frequency from 20MHz to 70MHz in steps of 1MHz, and record the outputs of (i) Digital frequency counter readout, and (ii) Phase tracker phase readout for the X and Y arms. I used the z avg -s utility to take an average for 10 seconds, and the standard deviation thus obtained correspond to the errorbars plotted. 
  5. Restore the connections to the green beat PDs and power them on again.

Y-arm transmission scan

I used the information from Attachment #2 to calibrate the X-axis of the Y-arm transmission data I collected on Wednesday evening. Looking at the beat frequency on the analyzer in the control room, between 24 and 47 MHz (green beat frequency, within the range the calibration was done over), we saw three IR resonances. I've marked these peaks, and also the 11MHz sideband resonances, in Attachment #3. It remains to fit the various peaks. I did a quick calculation of the FSR, and the number I got using these 3 peaks is 3.9703 +/- 0.0049 MHz. This value is ~23 kHz greater than that reported in elog 9804, but the error is also ~4 times greater (6 IR resonances were scanned in elog 9804) so I think these measurements are consistent.

Rubidium clock

I had brought an FS725 Rubidium clock back from W Bridge - the idea was to hook this up to the GPS 1PPS output, and use the 10MHz output from the FS725 as the reference for the fluke 6061A. However, the FS725 has not locked to the Rb frequency even though it has been left powered on for ~2days now. Do I have to do something else to get it to lock? The manual says that it should lock within 7 minutes of being powered on. Once this is locked, I can repeat the calibration with an 'absolute' frequency source...

Attachment 1: Xcalib.pdf
Xcalib.pdf
Attachment 2: Ycalib.pdf
Ycalib.pdf
Attachment 3: Y_scan_log.pdf
Y_scan_log.pdf
Attachment 4: 2015-11-05_phase_tracker_calib.dat.zip
Attachment 5: 2015-11-04_y_arm_scan.dat.zip
  11737   Thu Nov 5 10:48:21 2015 yutaroUpdateElectronicsoffset voltage vs. gain of common mode servo

I report the results of the measurement to know how offset voltage of common mode servo changes when the gain is changed.

 

- Motivation: If discontinuous change of the offset happens when we change the gain, it could cause saturation somewhere and so make the length control down. So, we want to estimate effect of such discontinuous change.

 

- Method: In 1 (or In 2) was terminated with 50 ohm, and the output voltage at Out 2 was measured with a multimeter (D040180-B).

 

- Results are shown below. Acquired data are attached in .zip.

The upper shows input equiv. offset. The lower shows offset measured at Out 2.

As for both In1 and In2, strange behaviors can be seen between -17 dB and -16 dB.

This is because 5 amplifiers (or attenuators) are simultaneously enabled/disabled here. Similar situation occurs every change of 8 dB gain. 

Attachment 2: offsets.zip
  11736   Thu Nov 5 03:04:13 2015 gautamUpdateCDSFrequency counting - systematics and further changes

I've made a few more changes to the frequency counting code - these are mostly details and the algorithm is essentially unchanged.

  • The C-code block in the simulink model now outputs the number of clock cycles (=1/16384 seconds) rather than the frequency itself. I've kept converting this period to frequency step by taking the reciprocal as the last step of the signal chain, i.e. after the LP filter.
  • In the current version of the FC library block, I've disabled the moving average in the custom C code - I've left the functionality in the code, but the window length at the moment is set to 1, which in effect means that there is no moving average. I found that comparable jitters in the output are obtained by using no moving average, and having two 2-pole IIR low-pass filters in series (at 4Hz and 2Hz) as by doing a moving average over 4096 clock cycles and then passing that through a 2Hz IIR lowpass filter (as is to be expected).

Systematic error

The other thing that came up in the meeting last week was this issue of the systematic errors in the measured frequency, and how it was always over-estimating the 'actual' frequency. I've been investigating the origin of this over the last few days, and think I've found an explanation. But first, Attachment #1 shows why there is a systematic error in the first place - because we are counting the period of the input signal in terms of clock cycles, which can only take on discrete, integer values, we expect this number to fluctuate between the two integers bounding the 'true value'. So, if I'm trying to measure an input signal of 3000Hz, I would measure its period as either 5 or 6 clock cycles, while the "true"  value should be 5.4613 clock cycles. In attachment #1, I've plotted the actual measured frequency and the measured frequency if we always undercounted/overcounted to the nearest integer clock cycles, as functions of the requested frequency. So the observed systematic error is consistent with what is to be expected.

The reason why this doesn't average out to zero is shown in Attachment #2. In order to investigate this further, I recorded some additional diagnostic variables. If I were to average the period (in terms of clock cycles - i.e. I look for the peaks in the blue cuve, add them up, and divide by the number of peaks), I find that I can recover the expected period in terms of clock cycles pretty accurately. However, the way the code is set up at the moment, the c code block outputs a value every 1/16384 seconds (red curve) - but this is only updated each time I detect a zero crossing - and as a result, if I average this, I am in effect performing some sort of weighted average that distorts the true ratio of the number of times each integer clock-cycle-period is observed. This is the origin on the systematic error, and is a function of the relative frequency each of the two integer values of the clock-cycle-period occurs, which explains why the systematic error was a function of the requested frequency as seen in Attachment #1, and not a constant offset. 

At the frequencies I investigated (10-70MHz in 5MHz steps), the maximum systematic error was ~1%.

Is there a fix?

I've been reading up a bit on the two approaches to frequency counting - direct and reciprocal. My algorithm is the latter, which is generally regarded as the more precise of the two. However, in both these approaches, there is a parameter known as the 'gate-time': this is effectively how long a frequency counter measures for before outputting a value. In the current approach, the gate time is effectively 1/16384 seconds. I would think that it is perhaps possible to eliminate the systematic error by setting the gate time to something like 0.25 seconds, and within the gate time, do an average of the total number of periods measured. Something like 0.25 seconds should be long enough that if, within the window, we do the averaging, and between windows, we hold the averaged value, the systematic error could be eliminated. I will give this a try tomorrow. This would be different from the moving average approach already explored in that within the gate-time, I would perform the average only using those datapoints where the 'running counter variable' shown in Attachment #2 is reset to zero - this way, I avoid the artificial weighting that is an artefact of spitting out a value every clock cycle. 

Attachment 1: Systematic_error.pdf
Systematic_error.pdf
Attachment 2: systematics_origin.pdf
systematics_origin.pdf
  11735   Thu Nov 5 02:18:32 2015 gautamUpdateLSCFSR and linewidth measurements with phase tracker

While the ETMx issues are being investigated - with Eric's help, I took some data from arm scans of the Y arm through ~2FSRs using ALS. I've also collected the data from the frequency counter readout during these scans but since they were done rather fast (over 60seconds), I am not sure how accurate this data will be. The idea however is to use the frequency readout from the phase tracker - this has to be linearized though, which I will do during the daytime tomorrow. The plan is to use our GPS timing unit to synchronize the following chain :

GPS timing unit 1PPS out --> FS725 Rb Clock 1PPS in (I recovered one which was borrowed from the 40m some time ago from the ATF lab yesterday evening, waiting for it to lock to the Rb clock now)

FS725 Rb Clock 10 MHz out --> Fluke 6061A 10MHz reference in

FS725 Rb Clock 10 MHz out-->agilent network analyzer 10MHz reference in (for measurement of the frequency of the signal output from the Fluke RF signal generator independent of its front panel display)

Then I plan to look at the phase tracker output as a function of the driving frequency (which will also be measured, offline, using the digital frequency counter setup) over a range of 20 MHz - 50 MHz in steps of 1 MHz. Results to follow.

Earlier tonight, Eric and I tweaked the PMC alignment (the mode cleaner was not staying locked for more than a couple of minutes, for almost an hour). 

  11733   Wed Nov 4 22:42:02 2015 ericqUpdateSUSETMX oplev servo disabled

During this time period, it looks like there was maybe one excursion. Here are the individual OSEM signals, which I think to be calibrated to microns. 

When I'm doing other things, I'm going to intentionally leave the watchdog tripped, and see if anything happens. 

  11732   Wed Nov 4 20:16:58 2015 yutaroUpdateElectronicsoffset voltage vs. gain of common mode servo

At 1Y2 rack, I measured offset voltage of the common mode servo (D040180-B) with the gain of it varied.

For now, all signal cables that come into or go out of the common mode servo are not plugged.

 

I will upload the data I took and report the result later.

  11731   Wed Nov 4 16:12:07 2015 ericqUpdateCDSOPTIMUS

There is a new machine on the martian network: 32 cores and 128GB of RAM. Probably this is more useful for intensive number crunching in MATLAB or whatever as opposed to IFO control. I've set up some of the LIGO analysis tools on it as well. 

A successor to Megatron, I dub it: OPTIMUS

  11730   Wed Nov 4 15:48:58 2015 SteveUpdatePEMETMX wirestadoff vs EQ

ETMX was not much effected by this 8.3 M Chilian earthquake. It was kicked up and it was kicking for hours after it. This is a sign of instability but it maintaind it's position.

The 24 hours plot follows the tempeerature.

 

Attachment 1: 8.3mChili.png
8.3mChili.png
  11729   Wed Nov 4 14:44:00 2015 ericqUpdateSUSETMX oplev servo disabled

I've turned the ETMX oplev servos off for the time being. (At the input side, so that no scripts will accidently turn it on).

Thus, only the local damping is being applied, let's see if we see any kicks...

  11728   Wed Nov 4 08:04:53 2015 SteveUpdatesafetysafety training

Yutaro Enamoto, visiting graduate student of Seiji received 40m specific basic safety training.

 

Attachment 1: Yutaro.jpg
Yutaro.jpg
  11727   Tue Nov 3 18:49:41 2015 gautamUpdateSUSETMX kicked up

I was trying to take a few more IR transmission scans with ALS when the ETMX got kicked again. I'm not sure how to fix this, so for the time being, I'm leaving the Oplev servo and the LSC turned off. The oplev spot looks really far off center especially in yaw, the yaw error is ~ -80.

Quote:

The oplev and  the LSC are off.

 

  11726   Tue Nov 3 03:12:46 2015 ericqUpdateLSCDRFPMI work

Tonight was kind of a wash. 

We spent some time retaking single arm scans with Gautam's frequency counting code to confirm the linewidths he measured before his most recent round of code improvements. During this, ETMX was being its old fussy self, costing us gentle realignment time. For the time being, we started actuating on ITMX for single arm locks. Also, out of superstition, I changed the static position offset that had been at +1k for the last N months to -1k. 

ETMX broke us out of a few DRFPMI lock trials as well, as did poor SRM alignment. I finally set up dither alignement settings for SRM in DRMI though, which helped (even in the arms-held-off-resonance situation). I still prefer doing the PRM/BS dither alignment in a carrier PRMI lock, because I think the SNR should be better than DRMI. 

We know that the ETMX excursions can happen without length drive exciting them, but also that length drives certainly can excite them. For future locks, I'm going to try out avoiding ETMX drive altogether; the sites use a single ETM for their DARM actuation and let the CARM loop take care of the resultant cross coupling, so hopefully we can do the same without angering the mode cleaner.

Anyways, we didn't really ever make it far enough to do anything interested with the DRFPMI tonight frown

  11725   Mon Nov 2 17:39:01 2015 KojiFrogsGeneralDRFPMI celebration

やったー!Yatta!

Attachment 1: yatta.jpg
yatta.jpg
  11724   Fri Oct 30 17:49:27 2015 SteveUpdateSUSETMX kicked up

The oplev and  the LSC are off.

Attachment 1: ETMXkicking.png
ETMXkicking.png
Attachment 2: ETMXkickedbyOpPerror.png
ETMXkickedbyOpPerror.png
Attachment 3: ETMXstillKicking.png
ETMXstillKicking.png
Attachment 4: ETMXkickingCond.png
ETMXkickingCond.png
Attachment 5: 7hrsETMX-Y.png
7hrsETMX-Y.png
  11723   Fri Oct 30 16:56:04 2015 gautamUpdateCDSis there a problem with the SCHMITTTRIGGER CDS library part?

Over the course of my investigations into the systematic errors in the frequency readout using digital frequency counting, I noticed that my counter variable that keeps track of the number of clock cycles between successive zero crossings was NOT oscillating between 2 values as I would have expected (because of there being a +/- 1 clock cycle difference between successive zero crossings due to the fixed sampling time of 1/16384 seconds), but that there were occassional excursions to values that were +/- 3 clock cycles away. I then checked the output of the SCHMITTTRIGGER CDS library part (which I was using to provide some noise immunity), and noticed that it wasn't triggering on every zero crossing at higher frequencies. I tested this by hooking up a digital oscillator to the SCHMITTTRIGGER part, and looked at its output for different frequencies. The parameters used were as follows:

CLKGAIN: 10000

SCHMITTTRIGGER lower threshold: -1.0

SCHMITTTRIGGER upper threshold: +1.0

I am attaching plots for two frequencies, 3000Hz and 4628Hz (Attachments #1 and #2) . I would have expected a flip in the state of the output of the schmitt trigger between every pair of horizontal red lines in this plot, but at 4628 Hz, it looks like the schmitt trigger isn't catching some of the zero crossings? Come to think of it, I am not even sure why the output of the schmitt trigger takes on any values other than 0 or 1 (could this be an artefact of some sort of interpolation in the visualization of these plots? But this would not affect the conclusion about the schmitt trigger missing some of the zero crossings?)

As an interim measure, I implemented a Schmitt trigger in my C code block - it was just a couple of extra lines anyways - I have designated the schmitt trigger output as a static variable that should hold its value in successive execution cycles, unless it is updated by comparing the input value to the thresholds (code attached for reference). Attachments #3 and #4 show the output of this implementation of a Schmitt trigger at the same two frequencies, and I am seeing the expected flip in the state between successive zero crossings as expected (though I'm still not sure why it takes on values other than 0 and 1?). Anyways this warrants further investigation. An elog regarding the implications of this on the systematic error in the frequency counter readout to follow.

Attachment 1: 3000Hz.pdf
3000Hz.pdf
Attachment 2: 4628Hz.pdf
4628Hz.pdf
Attachment 3: 3000Hz_software_SCHMITT.pdf
3000Hz_software_SCHMITT.pdf
Attachment 4: 4628Hz_software_SCHMITT.pdf
4628Hz_software_SCHMITT.pdf
  11722   Thu Oct 29 03:25:49 2015 ericqUpdateLSCDRFPMI work

[ericq, Gautam]

The length of DRFPMI lock did not increase much tonight, but we got a ~80 second sensing matrix measurement, and got the CARM bandwidth up to 10k with two boosts on. 

NB: I did not measure the CARM loop gain at its excitation frequency, so the plotted sensing element is supressed by the CARM loop. However, this is still useful for gauging the size of the PRCL signal vs. the residual CARM fluctuations. The excitations are fairly closely spaced between 309 and 316 Hz. 

For comparison, I'm also re-plotting the DRMI sensing measurement from a few weeks back taken at CARM offset of -4. We can see some change in the PRCL sensing, likely due to the CARM-coupled path. MICH/PRCL sadly looks pretty degenerate, but REFL55 looks more reasonable. 

 

I think the main limitation tonight was SRC stability. Even before bringing CARM to zero offset, we would see occasional sharp dives in AS110 power. One lockloss happened soon after such an occurance, but I checked the values, and it was not sufficient to trigger the Schmitt trigger down; instead it may have been a real optical loss of signal. The SRCL OLTF looks sensible. 

Random notes:

  • Aux X laser was glitching yet again, twiddled laser current to 1.90A from the 1.95A that I twiddled it to on Monday from the nominal 2.0A. 
  • When aligning the PRMI, I saw both ITMs' oplevs shift by a few urad in both pitch and yaw when engaing/breaking the lock, but this was not repeatable.
  • I reduced the AS110 whitening gain by 9dB, since the DC values were a few thousands, and I wanted to make sure there were no stray ADC saturations. This didn't change lock stability though. 
Attachment 1: DRFPMI.pdf
DRFPMI.pdf
Attachment 2: DRMIarms.pdf
DRMIarms.pdf
  11721   Wed Oct 28 17:06:30 2015 ericqUpdateASCNew PRC Angular FF filters installed

I've installed new filters for the T240 -> PRM static online angular feedforward that were trained after some of the recent changes to the signal chain of the relevant signals (i.e. the counts->velocity calibration that Rana did for the seismometers, and fixing the improper dewhitening of the POP QPD channels used as the Wiener target.)

Quickly trying them out now shows about the same level of performance as the previous ones, but the real performance I care about is during after-hours locking-time, so I'll take more measurements tonight to be posted here. 

  11720   Wed Oct 28 14:07:44 2015 KojiUpdateIOOMC WFS Offsets update

MC WFS offsets were updated to have a better operating point.
 

  11719   Tue Oct 27 10:08:12 2015 SteveUpdateVAC CC vacuum gauges

CC1 and CC2 are working again. Why did they start working  again ?

 

Attachment 1: gauges_5y_vs_8d.png
gauges_5y_vs_8d.png
Attachment 2: 13.9yCC1&4.png
13.9yCC1&4.png
  11718   Tue Oct 27 03:56:52 2015 ericqUpdateLSCDRFPMI work

A handful of DRFPMI locks tonight, longest one was ~7 minutes. 

EPICS/network latency has been a huge pain tonight. The locking script may hang between commands at an unstable place, or fail to execute commands altogether because it can't find the EPICS channel. This prevented or broke a number of locks. 

I made some CARM OLG and crossover measurements, and found the AO gain for the right crossover freq (~100Hz) to be ~8dB different than what's in the PRFPMI script, which is weird. Right now, the CARM bandwidth / ability to turn on boosts is limited by the gain peaking in the IMC CLG due to the high-ish PC/PZT crossover frequency we're using. 

Gautam turned on some sensing excitations during the last couple of locks, but they weren't on for very long before the lock loss. Hopefully I can pull out at least some angles from the data. 

I'm also more convinced that the PRC angular FF needs retuning; there is more residual motion on the cameras than I'm used to seeing. I've taken more data that I'll use to recalculate a wiener filter tomorrow. 

The PMC, ALSX beat and ITMX oplev all needed a reasonable pitch realignment tonight. 

  11716   Tue Oct 27 03:46:26 2015 ericqUpdateSUSMC2 F2P mis-tuned

D'oh. Good point.

Reverted for now; I'm thinking about doing laser pointer->MC2 QPD...

  11715   Mon Oct 26 19:10:59 2015 gautamUpdateGreen LockingAUX PDH loop characterization

I began my attempts to characterize the PDH loops at the X end today. My goal was to make the following measurements:

  • Dark noise and shot noise of the PD
  • Mixer noise
  • Servo electronics noise 

which I can then put into my simulink noise-budget scheme for the proposed IR beat setup.

I've made an Optickle model of a simple FP cavity and intend to match the measured PDH error signal from the X end to the simulated error signal to get the Hz/V calibration. I'll put the plots up for these shortly.

With regards to the other measurements, I was slowed down by remote data-acquisition from the SR785 - I've only managed to collect the analyzer noise floor data, and I plan to continue these measurements during the day tomorrow. 

  11714   Mon Oct 26 18:59:25 2015 gautamUpdateLSCGreen beatnote couplers installed

I found (an old) 10 dB coupler in the RF component shelves near MC2 - it has BNC connectors and not SMA connectors, but I thought it would be worth it to switch out the 20dB coupler currently on the X green beat PD on the PSL table with it. I used some BNC to SMA adaptors for this purpose. It appears that the coupler works, because I am now able to register an input signal on the X arm channel of the digital frequency counter (i.e. the coupled output from the green beat PD). I thought it may be useful to have this in place and do an IR transmissions arm scan using ALS for the X arm as well, in order to compare the results with those discussed here. However, the beat note amplitude on the analyzer in the control room looks noticeably lower - I am not sure if the coupler is responsible for this or if it has to do with the problems we have been having with the X end laser (the green transmission doesn't look glitchy on striptool though, and the transmission itself is ~0.45). In any case, we could always remove the coupler if this is hindering locking efforts tonight. 

  11713   Mon Oct 26 18:10:38 2015 IgnacioUpdateIOOLast Wiener MCL subtractions

As per Eric's request, here is the code and TF measurement that was used to calculate the MC2 FF filter that is loaded in FM5. This filter module has the filter with the best subtraction performance that was achieved for MCL.

code_TF.zip

Attachment 1: code_TF.zip
  11712   Sat Oct 24 12:34:43 2015 gautamUpdateCDSFrequency counting - workable setup prepared

Sorry for the confusion - I did mean Green beat frequency, and I had neglected the factor of 2 in my earlier calculations. However, the fit parameter "c" in my fit was actually the half-width at half maximum and not the full width at half maximum. After correcting for both these errors (new fit is Attachment #1, where I have now accounted for the factor of 2, and the X axis is the IR beat frequency), I don't think the numbers change too much. It could be that the frequency counter wasn't reading out the frequency correctly, but looking at a time series plot of the frequency counter readout (Attachment #2), and my earlier trials, I don't think this is the case (38 MHz is a frequency at which I don't expect much systematic error - also, the offset was stepped from -3 to 3 over 45 seconds). 

The revised numbers:

Fitted linewidth = 2*c = 10.884 kHz +/- 2 Hz (95% C.I.)

FSR of Y arm (from elog 9804) = 3.9468 MHz +/- 1.1 kHz

=> Y arm Finesse = FSR/fitted linewidth =  362.6 +/- 0.5

Total round trip loss = 2*pi/Finesse = 0.0173

 

Attachment 1: Yscan.pdf
Yscan.pdf
Attachment 2: Frequency_readout.pdf
Frequency_readout.pdf
  11711   Fri Oct 23 21:58:10 2015 ranaUpdateSUSMC2 F2P mis-tuned

The OSEMs cannot be used for coil balancing above ~10 Hz. The main coupling path from OSEM drive to sensor is not through the mirror motion, but instead direct electrical coupling of the drive wires to the sensing wires.sad

I put this in the elog every ~1-2 years since people keep trying it, but it keeps coming back like a zombie.crying

Better to use the MC angular sensors for L2A decoupling. Not perfect, but better than OSEMs. For the TMs we can use the OLs.

  11710   Fri Oct 23 19:27:19 2015 KojiUpdateCDSFrequency counting - workable setup prepared

It is a nice scan. I'm still thinking about the equivalence of the moving average and the FIR low pass that I have mentioned in the meeting.

Anyways:

I'm confused by the plot. The bottom axis says "green beat frequency".
If you scan the IR laser frequency by df, you get 2*df shift of the green beatnote. You need to have this factor of two somewhere.
If you are looking at the IR beat freq, just the label is not correct. (I believe this is the case.)

Accepting the rather-too-low finesse of 363 (nominally 450), the total round trip loss is said to be 0.0173.
If we subtract the front transmission of 1.38% (and ignoring the transmission loss from the ETM),
the round trip loss is 3500ppm. Is this compatible with the following elog?
http://nodus.ligo.caltech.edu:8080/40m/11111
In fact, I'm afraid that the loss number in the above elog 11111 was not correct by a factor of 10.
Then, if so, can we believe this high loss number? (Nominally we expect ~100ppm loss per round trip...)

  11709   Fri Oct 23 18:36:48 2015 gautamUpdateCDSFrequency counting - workable setup prepared

I've made quite a few changes in the software as well as the hardware of the digital frequency counting setup.

  • The main change on the software side is that the custom C code that counts intervals between successive zero crossings and updates the frequency output now has a moving average capability - the window size is readily changable (by a macro in the first line of the code, which resides at /opt/rtcds/userapps/release/cds/c1/src/countZeroCrossingWindowed.c - however, changing the window size requires that the model be recompiled and restarted), and is currently set to 4096 because based on some empirical trials I did, this seemed to give me the frequency output with the least jitter, and also smaller systematic errors than in my earlier trials described here.
  • The filter modules for both the X and Y channels now have 2 pole butterworth low pass filters with poles at 64Hz, 32Hz, 16Hz, 8Hz, 4Hz, 2Hz and 1Hz loaded. Again, based on my empirical trials, a combination of a moving average filter in the C code and the IIR filters after that worked best in terms of reducing the jitter in the frequency readout. I think the fact that the moving average 'spreads' the impulse caused by a glitch in the counting algorithm improves the response of the combination as compared to having only the IIR filters in place. 
  • The Frequency Counting SIMULINK block has been cleaned up a little - I have removed unnecessary test points I had set up for debugging, and is now a library part called "FC".
  • After the experience of having C1ALS crash as noted here, I was doing all my testing in the C1TST model. Having made all the changes above, I reverted to the C1ALS model, which compiled and ran successfully this time.
  • On the hardware side, I interchanged the couplers mentioned here - so the 20dB coupler now sits on the X green beat PD, while the 10dB coupler sits on the Y green beat PD. This change was motivated by wanting to test out the digital frequency counting setup by performing an arm scan through an IR resonance using ALS, and we found that the PSL+Y green beat frequency was better behaved than the X+PSL combination.

Arm scan

Eric helped me test the new setup by doing an arm scan through an IR resonance by ramping the ALS offset from -3 to +3 with a ramp time of 45 seconds. The data was acquired with the window size of the moving average set to 4096 clock cycles, and a 2 Hz low pass IIR filter before the frequency readout. Attachment 1 shows a plot of the data, and a fit with a function of the form trans = a/(1+((x-b)/c)^2), where a = normalization, b = center of lorentzian, and c = linewidth (FWHM) of the peak (the fitted parameter values, along with 95% confidence bounds are also quoted on the plot). In terms of the data acquisition, comparing this dataset to one from an earlier scan Eric did (elog11111) suggests that the frequency counting setup is working reasonably well - at any rate, I think the data is a lot cleaner than before implementing the moving average and having a 20Hz lowpass IIR filter. In any case, we plan to repeat this measurement sometime next week during a nighttime locking session. It remains to calculate the arm loss from these numbers analogous to what was done earlier for the X arm.

Calculation of loss:

Fitted linewidth = 10.884 kHz +/- 11Hz (95% C.I.)

FSR of Y arm (from elog 9804) = 3.9468 MHz +/- 1.1 kHz

=> Y arm Finesse = FSR/fitted linewidth =  362.6 +/- 0.5

Total round trip loss = 2*pi/Finesse = 0.0173

 

 

 

 

Attachment 1: Yscan.pdf
Yscan.pdf
  11708   Fri Oct 23 09:55:50 2015 SteveUpdateLSCstable days
Attachment 1: stable4days.png
stable4days.png
  11707   Thu Oct 22 08:52:04 2015 SteveUpdateVACclean RGA scan after sweaty Maglev

Clean comparable scan at vacuum normal. There was no backstreaming.

Attachment 1: clean_scan.png
clean_scan.png
  11706   Wed Oct 21 15:42:56 2015 SteveUpdateVACvacuum gauge check

Convectron gauge check:

Brand new 10 years old convectron gauge at atm was swapped into the place of existing gauges to see if they read close to 760 Torrs

They did reasonable well at the low end. I tried to imitate calibration and bring down the high end with little success.

( P1 and P2  were reading 7e-4 to ~660 Torr  The correction of the upper end pushed up the lower end too. I will correct this later

P3 and P4 high ends are way off )

The point is that they work.  Convectron gauges will be replaced and calibrated at the next vent.

Interlocks were not triggered during this test. I was expected to close the PSL shutter when P1 was reading 760 Torr

This hide some problem or  not understanding.

It was good to see CC1 and CC4 working at the moment

 

Somehow I succeded opening VM1 and closed VM2 = vacuum normal

I just hope it stays open overnight to get comparable RGA scan.

 

Attachment 1: Gauges.png
Gauges.png
  11705   Wed Oct 21 10:03:15 2015 SteveUpdateVACafter running out of N2

 

Quote:

[ericq, Gautam, Steve]

Following roughly the same procedure as ELOG 11354, c1vac1 and c1vac2 were rebooted. The symptoms were identical to the situation in that ELOG; c1vac1 could be pinged and telneted to, but c1vac2 was totally unresponsive. 

The only change in the linked procedure was that we did not shut down the maglev. Since I unwittingly had it running for days without V4 open while Steve was away, we now know that it can handle shorter periods of time than that...

Upon reboot, many channels were readable again, unfortunately the channels for TP2 and TP3 are still blank. We were able to return to "Vacuum normal state," but because of unknowned communication problems with VM1's interlock, we can't open VM1 for the RGA. Instead we opened VM2 to expose the RGA to the main IFO volumn, but this isn't part of the "Normal" state definite, so things currently read "Undefined state".

Precondition:

1, Pressure gauges had no communication ( NO COMM ) with c1vac2

 

2, Lost N2 supply on Oct 9 This triggered a normal all valve closed condition. At this point you replace N2 cylinders and manually swich valves to recreate VAC NORMAL configuration in the correct sequential order.

   The very last thing you do is open V1 gate valve.

   a, check TP2 that is the forepump of the Maglev. Foreline pressure to drypump  ~ 10- 100 mTorr, rotation speed 50 Krpm

   b, open V4 if P2 <1Torr

   c, check Maglev rotating at 560 Hz

   d, open V1 if P1 <500 mTorr

   e, check TP3 foreline, rotation speed and open V5 if P3 <1 Torr with VA6 closed

   f, open VA6 if PAN <1 Torr

  g, open annulos valves one by one , like VASE if PASE <1 Torr and so on...........Now the Current State: should read Vac Normal

 

3, Maglev run for 7 days with V4 closed. This encreased its foreline pressure to estimated few Torrs  and its body temp rose ~30C on the outside.

   So it was sweating and it may be back streamed.

The present RGA data is indicating that it had to be very mild.

The RGA will have better sensitivity with VM1 open and VM2 closed.

The PSL output shutter stayed open during these period is pointing  to IFO pressure stayed P1 <3 mTorr

PROBLEM: P1 and P2 plot should show nothing where there is no communication.http://nodus.ligo.caltech.edu:8080/40m/151016_182003/oct16Fpm2015.png

                   How do we check if pressure based software interlocks are working in this no communication condition?

Attachment 1: 1d_after_reboot.png
1d_after_reboot.png
  11704   Tue Oct 20 17:36:01 2015 gautamUpdateCDSFrequency counting with moving average

I'm working on setting up a moving-average in the custom C code block that counts the zero crossings to see if this approach is able to mitigate the glitchy frequency readout due to mis-counting by one clock cycle between successive zero crossings. I'm storing an array the size of the moving average window of frequency readouts at each clock cycle, and then taking the arithmetic mean over the window. By keeping a summing variable that updates itself each clock cycle, the actual moving average process isnt very intensive in terms of computational time. The array does take up some memory, but even if I perform the moving average over 1 second with 16384 double precision numbers stored in the array, its still only 130 kB so I don't think it is a concern. Some tests I've been doing while tuning the code suggest that with a moving average over 16384 samples (i.e. 1 second), I can eliminate glitches at the 1Hz level in the frequency readout for frequencies up to 5 kHz (generated digitally using an oscillator block). Some tuning still needs to be done, and the window could possibly be shortened. I also need to take a look at the systematic errors in this revised counting scheme, preferably with an analog source, but this is overall, I think, an improvement.

On a side note, I noticed some strange behaviour while running the cds average command - even though my signal had zero fluctuations, using z avg 10 -s C1:TST-FC_FREQUENCY_OUT gave me a standard deviation of ~1 kHz. I'm not sure what the problem is here, but all the calibration data I took in earlier trials were obtained using this so it would be useful to perform the calibration again. 

  11703   Tue Oct 20 16:27:04 2015 ericqUpdateVACvacuum VME machines rebooted

[ericq, Gautam, Steve]

Following roughly the same procedure as ELOG 11354, c1vac1 and c1vac2 were rebooted. The symptoms were identical to the situation in that ELOG; c1vac1 could be pinged and telneted to, but c1vac2 was totally unresponsive. 

The only change in the linked procedure was that we did not shut down the maglev. Since I unwittingly had it running for days without V4 open while Steve was away, we now know that it can handle shorter periods of time than that...

Upon reboot, many channels were readable again, unfortunately the channels for TP2 and TP3 are still blank. We were able to return to "Vacuum normal state," but because of unknowned communication problems with VM1's interlock, we can't open VM1 for the RGA. Instead we opened VM2 to expose the RGA to the main IFO volumn, but this isn't part of the "Normal" state definite, so things currently read "Undefined state".

  11702   Tue Oct 20 15:51:44 2015 ericqUpdateSUSMC2 F2P tuned

Using a modified version of Hang's deMod_deCoup scripts, I tuned the MC2 coil output matrix to minimize the appearance of POS drive in the SUSPIT signal at 28Hz. Up until now, there was no F2P compensation. This reduced the force to pitch coupling at 28Hz by 8dByes

Old: POS -> 1 x UL, 1 x UR, 1 x LL, 1x LR

New: POS -> 1.1054 x UL, 1.1054 x UR, 0.8946 x LL, 0.8946 LR

I checked the MCL spectrum before and after this change with OAF on, this did not spoil the feedforward length subtraction in any noticible way. 

The script lives in userapps/release/isc/c1/scripts/decoup, but I've symlinked it to /opt/rtcds/caltech/c1/scripts/decoup.


The script modification I made had to do with how the I and Q data is collected. Before, it was sporadically probing the I and Q FM output monitor EPICS values; I changed it to use the avg function of cdsutils, which calculates the mean and std from the 16kHz data and have seen it improve results by 1dB or so. I've been in touch with Jenne to propagate this to the sites. 

  11701   Tue Oct 20 11:24:29 2015 ericqHowToLSCHow to DRFPMI

Herein, I will describe the current settings and procedures used to achieve the DRFPMI lock, cobbled together from scripts, burts and such. 


Initial Alignment

  1. With arms POX/POY locked, run dither alignment servos. Set transmon QPD offsets here
  2. Restore "PRMI Carrier" configuration, run BS and PRM dither alignment servos simultaneously. (Note: this sacrifices some X arm alignment for better dark port alignment. In practice no appreciable loss of TRX is observed)
  3. Misalign PRM, align SRM and tune SRM alignment by eye while looking at AS camera. 
  4. Restore POX/POY arm lock, lock green to arms, check that powers are high enough and align if neccesary.

Initial Configuration

CARM, DARM

For CARM and DARM, the A channels are used for the ALS signals, whereas the B channels are used for blending the RF signals. 

ALS

  • BEATX and BEATY, I and Q channels: +0dB Whitening Gain, Whitening Filters ON
  • Green beatnotes somewhere between 20-80MHz, following sign convention of temperature slider UP makes beat freq go UP.  Check spectrum of PHASE_OUT_HZ vs references in ALS_outOfLoop_Ref.xml. The locking script automatically sets the correct phase tracker gain, so no need to adjust manually.
  • CARM_A = -1.0 x ALSX + 1.0 x ALSY, G=1.0
  • DARM_A = 1.0 x ALSX + 1.0 x ALSY, G=1.0

RF

  • CM Board: REFL11 I daugher board output -> IN1, IN1 Enabled, -32dB input gain, 0.0V offset, all boosts off, AO polarity positive, AO gain +0dB
  • MC Board: IN2 disabled, -32dB input gain
  • CM_SLOW: +0dB Whitening Gain, Whitening ON, LSC-CM_SLOW_GAIN = -5e-4 (Though, it would be good to reallocate this gain to the input matrix element)
  • CARM_B = 1.0 x CM_SLOW, FM4 FM10 ON, G=0 (FM4 = LP700 for AO crossover stability, FM10 = 120:5k for coupled cavity pole compensation)
  • AS55: +9dB Whitening Gain, Whitening filters manual, Demod angle -37.0
  • DARM_B = -1e-4 x AS55 Q, G=0

DRMI 3F

For the DRMI, the A channels are used for the 1F signals, whereas the B channels are used for the 3F signals. The settings for transitioning to 1F after locking the DRFPMI have not yet been determined. 

These settings are currently saved in the DRMI configurator, but the demod angles are set for DRFPMI lock, so the settings don't reliably work for misaligned arms. 

  • REFL33: +30dB Whitening Gain, Whitening filters trigger on DRMI lock, Demod angle: 136.0
  • REFL165: +24dB Whitening Gain, Whitening filters trigger on DRMI lock, Demod angle: -111.0
  • POP22: +15dB Whitening Gain, Whitening filters OFF, Demod angle: -114.0
  • AS110: +36dB Whitening Gain, Whitening filters OFF, Demod angle: -116.0
  • POPDC: +0dB Whitening Gain, Whitening filters OFF (used as a supplemental trigger signal when CARM and DARM are buzzing and POP22 fluctuates wildly)
  • MICH_B = 6.0 x REFL165Q, offset = 15.0
  • PRCL_B = 5.0 x REFL33I, offset = 45.0
  • SRCL_B = -0.6 x REFL165I + 0.24 x REFL33 I, offset=0

The REFL33 element in SRCL_B is to reduce the PRCL coupling, was found empirically by tuning the relative gains with the arms misaligned and looking at excitation line heights. The offsets were found by locking the DRMI on 1F signals with arms misaligned, and taking the average value of these 3F error signals.  

Servo filter configuration

The CARM and DARM ALS settings are largely scripted by scripts/ALS/Transition_IR_ALS.py, which takes you from arms POX/POY locked to CARM and DARM ALS locked. The DRMI settings are usually restored from the IFO_CONFIGURE screen. 

  • CARM: FM[1, 2, 3, 5, 6] , G=4.5, Trigger forced on, no FM triggers, output limit 8k
  • DARM: FM[1, 2, 3, 5, 6] , G=4.5, Trigger forced on, no FM triggers, output limit 8k
  • MICH: FM[4, 5], G= -0.03, Trigger POP22 I x 1.0 [50, 10], FM[2, 3, 7] triggered [50, 10], output limit 20k
  • PRCL: FM[4, 5], G= -0.003, Trigger POP22 I x 1.0 [50, 10], FM[1, 2, 8, 9] triggered [50, 10], output limit 8k
  • SRCL: FM[4, 5], G= -0.4, Trigger AS110 Q x 1.0 [500, 100], FM[2, 7, 9] triggered [500, 100], output limit 15k

Actuation Output matrix

  • MC2 = -1.0 x CARM
  • ETMX = -1.0 x DARM
  • ETMY = 1.0 x DARM
  • BS = 0.5 x MICH
  • PRM = 1.0 x PRCL - 0.2655 MICH
  • SRM = 1.0 x SRCL + 0.25 MICH (The mich compensation is very roughly estimated)

Locking Procedure

When arms are POX/POY locked, and the green beatnotes are appropriately configured, calling scripts/DRFPMI/carm_cm_up.sh initiates the following sequence of events:

  • Turn ON MC length feedforward and PRC angle feedforward
  • Set ALS phase tracker UGFs by looking at I and Q magnitudes
  • Set LSC-ALSX and LSC-ALSY offsets by averaging, ramp CARM+DARM gains up, XARM+YARM gains down, engage CARM+DARM boosts, now ALS locked
  • Move CARM away from resonance, offset = -4.0 (DRMI locks quicker on this side for whatever reason)
  • Restore PRM, SRM alignment. Set DRMI A FM gains to 0, B FM gains to 1.0. Enable LSC outputs for BS, PRM, SRM
  • When DRMI has locked, add POPDC trigger elements to DRMI signals and transition SRCL triggering to POP22I. NB: In the c1lsc model, the POPDC signal incident on the trigger matrix has an abs() operator applied to it first. 
    • MICH Trig = 1.0 x POP22 I + 0.5 x POPDC, [50, 10]
    • PRCL Trig = 1.0 x POP22 I + 0.5 x POPDC, [50, 10]
    • SRCL Trig = 10.0 x POP22 I + 5 x POPDC, [500, 100]
  • Reduce POX, POY whitening gains from their nominal +45dB to +0dB, so there aren't railing channels making noise in the whitening chassis and ADCs
  • DC couple ITM oplevs (average spot position, set FM offset, turn on DC boost filter, let settle)
  • With an 8 second ramp, reduce CARM offset to 0 counts. 
  • MANUALLY adjust CARM_A and DARM_A offsets to where CARM_B_IN and DARM_B_IN are seen to fluctuate symetrically around their zero crossing.
    • Note: Last week, this adjustment tended to be roughly the same from lock to lock, unlike the PRFPMI which generally didn't need much adjustment. Also, by jumping from CARM offset of -0.4 to 0.4, it could be seen that the zero crossing in  CARM_B aka CM_SLOW aka REFL11 had some offset, so CARM_B_OFFSET was set to 0.005, but this may change. 

When CARM and DARM are buzzing around true zero, powers maximized:

  • CARM and DARM FM1 (18,18:1,1 boosts) OFF
  • CARM_B_GAIN 0.0 -> 1.0, FM7 ON (20:0 boost)
  • DARM_B_GAIN 0.0 -> 0.015, FM7 ON (20:0 boost) 
  • MC servo board IN2 ENABLE, IN2 gain -32dB -> -16dB
  • Turn ALL MC2 violin filters OFF (smoothen out AO crossover)
  • If stable, CM board IN1 gain -32dB -> -10dB (This is the overall CARM gain, the arm powers stabilize within the last few dB of this transition)
  • CARM_A_GAIN 1.0 -> 0.7
  • CARM_A FM9 ON (LP1k), sleep, FM 1 ON (1:20 deboost), sleep, FM 2 ON (1:20 deboost), HOLD OUPUT, CARM now RF only
  • DARM_B_GAIN 0.015 -> 0.02, sleep, DARM_A_GAIN 1.0 -> 0.0 (This may not be the ideal final DARM_B gain, UGF hasn't been checked yet)

IFO is now RF only!

  • Turn on transmon QPD servos.
  • Adjust comm/diff QPD servo offsets to correct any problems evident on AS/REFL cameras. This usually brings powers from ~100-120 to ~130-140. 

This is as far as we've taken the DRFPMI so far, but the CARM bandwidth is still only at a few kHz. Based on PRFPMI locking, the next steps will be:

  • CM BOARD +12dB or so additional IN1 gain, more AO gain may be needed to get crossover to final position of ~100Hz
  • MC2 viollin filters back on
  • CM boost(s) on
  • AS55 whitening on
  • Transition DRMI to 1F
  11700   Mon Oct 19 16:20:52 2015 ericqUpdateLSCGreen beatnote couplers installed

Last Friday, I installed some RF couplers on the green BBPDs' outputs, and sent them over to Gautam's frequency divider module. At first I tried 20dB couplers, but it seemed like not enough power was reaching the dividers to produce a good output. I could only find one 10dB coupler, and I stuck that on the X BBPD. With that, I could see some real signals come into the digital system.

I don't think it should be a problem to leave the couplers there during other activities.

  11699   Mon Oct 19 16:16:01 2015 ericqUpdateCDSC1ALS crashes

Gautam was working on his digital frequency counter stuff when the c1als model crashed. I had trouble bringing it back until I realized that, for reasons unknown to me, the safe.snap file that the model looks for at boot had been deleted. (This file lives in /target/c1als/c1alsepics/burt). I copied over this morning's version from Chiara's local backup. 

At the sites, these files are under version control in the userapps svn repository, presumably symlinked into the target directory. We should definitely do something along these lines. 

  11698   Mon Oct 19 15:23:22 2015 ericqUpdateLSCLonger DRFPMI lock

Here is a longer lock, about 100 seconds RF only, from later that same night. The in-loop CARM and DARM error signals have the order of magnitude of 1nm per count. 

From ~-150 to -103, we were fine tuning the ALS offsets to try and get close to the real CARM/DARM zero points then blending the RF CARM signal. 

At -100, the CARM bandwidth increases to a few kHz and stabilizes the arm powers. By -81, the error signals are all RF. At -70, I turned on the transmon QPD servos, which brought the power up a bit. 

If I recall correctly, lock was lost because I put waaaay too big of an excitation on DARM with the goal of running its UGF servo for a bit. The number I entered was appropriate for ALS, but most certainly too huge for AS55...

  11697   Mon Oct 19 11:20:34 2015 gautamUpdateVACRGA scan reset

Steve pointed out that in the aftermath of the Nitrogen running out a couple of times last week, the RGA had shut itself off thinking that there was a leak and so it was not performing the scheduled scans once a day. So the data files from the scheduled scans were empty in the /opt/rtcds/caltech/c1/scripts/RGA/logs directory. The wiki page for getting it up and running again is up-to-date, but the script RGAset.py did not exist on the c0rga machine, which the RGA is communicating with via serial port. I copied over the script RGAset.py from rossa to c0rga and ran the script on that machine - but the error flags it returned were not all 0 (indicating some error according to the manual) - so I edited the script to send just the initialize command ('IN0') and commented out the other commands, after which I got error flags which were all 0. After this, I ran a manual scan using 'RGAlogger.py', and it appears that the RGA is now able to take scans again - I'm attaching a plot of the scan results. We've saved this scan as a reference to compare against after a few days. 

Attachment 1: RGAscan_151019.png
RGAscan_151019.png
  11696   Sat Oct 17 18:55:07 2015 ranaUpdateLSCDRFPMI Locked for 20 sec

smiley in addition to Koji's words I feel like we should also thank those who made small but positive contributions. Its hard not to notice that this locking only happened after the new StripTool PEM colors were implemented...

From the times series plot I guess that the fuzz of the in-loop DARM is 1 pm RMS (based on memory). This means that the ALS was holding the DARM at 10 pm from the RF resonances.

There is no significant shift in the DRMI error signals, so new weird CARM effect. Would be interesting to see what the 1f signals do in the last 60 seconds before RF lock.

For documentation, perhaps Gautam can post the loop gain measurements of the 5 loops on top of the Bode plots of the loop models.

  11695   Fri Oct 16 18:27:52 2015 SteveUpdateVAC vacuum normal condition

Vacuum normal configuration with VM1-closed, VM2-open valve positions. Power load normal 24V 0.2A

Maglev rotation 560 Hz at room temp body temp.

"NO COMM" error message on medm screen. Gauge controller pressures are read able.

All vac comp LEDs are green. We have to reboot on Monday to enable communication.

 

Attachment 1: oct16Fpm2015.png
oct16Fpm2015.png
  11694   Thu Oct 15 14:39:58 2015 ericqUpdateComputer Scripts / Programsnodus web apache simlinks too soft
Quote:

None of the links here seem to work. I forgot what the story is with our special apache redirect frown

https://wiki-40m.ligo.caltech.edu/Core_Optics

The story is: we currently don't expose the whole /users/public_html folder. Instead, we are symlinking the folders from public_html to /export/home/ on nodus, which is where apache looks for things

So, I fixed the links on the Core Optics page by running:

controls@nodus|~ > ln -sfn /users/public_html/40m_phasemap /export/home/

  11693   Thu Oct 15 10:59:12 2015 KojiUpdateLSCDRFPMI Locked for 20 sec

Great job!
Many thanks Eric, Gautam, and all the current and past colleagues
for your tremendous contributions to bring the 40m to this achievement.

  11692   Thu Oct 15 04:14:14 2015 ericqUpdateLSCDRFPMI Locked for 20 sec

Fast ALS was still a problem tonight. I don't think high frequency ALS noise saturating the PC drive is the issue; I put two 10k poles before the CM board (shooting for just 2-3kHz bandwidth), and the PC drive levels would be stable and low up until the lockloss, which was always conincident with a step in the AO gain.

After working with that for a few hours, we turned back to our more standard locking attempts. First, we dither aligned the PRMI, and then centered the REFL beam on REFL11. It's hard to say for certain, but we may have been a little close to the edge of the PD. The only other thing that differed from Monday's attempts was using 6dB less AO gain when trying the up the overall gain. 

The script now reliably breaks through to stable high powers, we had a handful of pure-RF locks tonight. The digital DARM gain needs tuning, and the CARM bandwidth still isn't at its final state, but these are very tractable. Off the top of my head, the way forward now includes:

  • Set proper final DARM loop shape
  • Set final CARM loop shape
  • Take full sensing matrix
  • Make 1F handoff
  • Set up the CAL model to produce (at least roughly) calibrated spectra
  • measure noise couplings and other fun stuff

Unrelated: I feel that the PRC angular FF may have deteriorated a bit. I'm leaving the PRC locked on carrier to collect data for wiener filter recalculation. 

  11691   Thu Oct 15 03:08:57 2015 ericqUpdateLSCDRFPMI Locked for 20 sec

[ericq, Gautam]

For real this time.

Attachment 1: DRFPMI_locked.pdf
DRFPMI_locked.pdf
  11690   Wed Oct 14 17:40:50 2015 gautamUpdateCDSFrequency divider box - further tests

Summary:

I carried out some further diagnostics and found some ways in which I could optimize the zero-crossing-counting algorithm, such that the error in the measured frequency is now entirely within the expected range (due to a +-1 clock cycle error in the counting). We can now determine frequencies up to ~60 MHz with less than 1 MHz systematic error and <10 kHz statistical error (fluctuations after the 20 Hz lowpass). This should be sufficient for slow control of the end-laser temperatures.

Details: 

The conclusion from my earleir tests was that there was possibly an improvement that could be made to setting the thresholds for the Schmitt trigger stage in the model. In order to investigate this, I wanted to have a look at the 64K sampled raw input to the ADCs. Yesterday Eric helped me edit the appropriate .par file for viewing these channels for c1x03, and for an input frequency of 70MHz (after division, ~4.3 kHz square wave), the signal looked as expected (top left plot, attachment #1). This prompted me to check the counting algorithm again with the help of various test points I had setup within the model. I found that there was a tendency to under-count the number of clock-cycles between zero-crossings by more than 1 clock cycle, due to the way my code was organized. I fixed this and found that the performance improved dramatically, compared to my previous trials. With the revised counting algortihm, there was at most a +-1 clock cycle error in the counting, and the systematic error between the measured and requested RF frequencies is now completely accounted for taking this consideration into account. The origin of this residual error can be understood by looking at the top right plot in Attachment #1 - presumably because of the effects of some downsampling filter, the input signal to the Schmitt trigger isnt a clean square wave (even at 4kHz) - specifically, the time spent in the LOW and HIGH states of the Schmitt trigger can vary between successive zero crossings because of the shape of the input waveform. As a result, there can be a +-1 clock cycle error in the counting process. Attachment #2 shows this - the red and blue lines envelope the measured frequency for the whole range investigated: 10-70MHz. Attachment #3 shows the systematic error as a function of the requested frequency.

If there was some way to bypass the downsampling filter, perhaps the high-frequency performance could be improved a little. 

Attachment 1: time_series_input_signals.pdf
time_series_input_signals.pdf
Attachment 2: calibration_20151012.pdf
calibration_20151012.pdf
Attachment 3: systematics_20151012.pdf
systematics_20151012.pdf
  11689   Wed Oct 14 16:53:22 2015 KojiUpdateGeneralNon inverting buffer for SR560

Eric needed a buffer to drive low input impedance (~130Ohm) of his pomona box, I quickly made a non-iverting buffer with G=+10. The DC power is obtained from the back of SR560. It uses 1.02K and 9.09K
to have the gain of ~10. The chip is OP27. In fact this limits the output swing to be +/-5V for the load resistance of 130Ohm. Eric thinks this is enough. If we need more, we need to swap the chip.

As SR560s tend to saturate too quickly, it would be very useful to have this kind of kit in all the labs
once it is packed in a box.

Attachment 1: IMG_20151014_145608547.jpg
IMG_20151014_145608547.jpg
Attachment 2: IMG_20151014_145548751.jpg
IMG_20151014_145548751.jpg
  11688   Wed Oct 14 15:59:06 2015 ranaUpdateComputer Scripts / Programsnodus web apache simlinks too soft

None of the links here seem to work. I forgot what the story is with our special apache redirect frown

https://wiki-40m.ligo.caltech.edu/Core_Optics

ELOG V3.1.3-